You need to score at least 1500 points on your new video game to obtain the high

score. You get 300 points after completing each level. Write and solve an inequality to

find the number of levels you must

beat in order to obtain the high score.


1
Expert's answer
2021-02-24T06:50:24-0500

Let "x=" the number of levels you must beat in order to obtain the high score.

The we have the inequality


"300x\geq1500"

Solve for "x"


"\dfrac{300x}{300}\geq\dfrac{1500}{300}"

"x\geq5"

You must beat at least 5 numbers in order to obtain the high score.